Fred is a 56 year old, single male. He is employed in sales and spends a great portion of his day either behind his desk making calls or in his car. Fred has a base salary but depends on sales for commission. He has a limited income and attempts to use finances wisely. Due to his schedule, he has few friends and little leisure time. He relays that he has been diagnosed with high blood pressure and has a family history of cardiovascular disease.

In a typical day, Fred sets his alarm for 7 am and tries to get to the office by 8:30. He always drives through a donut shop and picks up a reduced fat blueberry muffin and a large latte.

For lunch he packs a bologna sandwich, a bag of Corn chips and a Cola and eats at his desk.

After making calls and finalizing several deals, he leaves the office around 6 pm. Several times a week, he meets clients for dinner. Today he is having dinner with a couple of guys at a local steak house.

After dinner and a business session, he returns home around 10 pm. Fred is exhausted but decides to unwind and watch the nightly news before bed.

Fred’s company is collaborating with an insurance company to begin a wellness program. As the nurse for the wellness program, you are to meet with each employee and complete an initial evaluation.

Answer the following questions pertaining to Fred’s evaluation.

Use myplate.gov (Links to an external site.)Links to an external site to answer the questions.

Client: Fred

Age: 56

Height: 5 feet 10 inches                            Weight: 270 pounds         

1. Daily calorie allowance for the day per myplate plan for Fred’s Height, weight and age_____________

2. Calories consumed: _____________________

3. % of calories from fat : ____________

4. Recommended maximum % of calories from fat:   ___________________

5. Based on the recommended calorie allowance, and using the MyPlate Plan, how many servings of each food group should Fred have daily?

Fruit:

Vegetables:

Grains:

Protein:

Dairy:

6. Looking at Fred’s diet recall and the MyPlate recommendations:

a. What food groups are inadequate in his diet?

b. What are the three concepts that MyPlate emphasizes to encourage healthy eating habit?

c. The Dietary Guidelines can be used by consumers to help them make changes to their eating habits that will last for a lifetime. Review the Dietary Guidelines MyPlate messages. Explain the 5 messages to Fred.

7. What information has Fred provided concerning his environmental assessment?

8. Before concluding this session, you provide recommendations for a healthier lunch. Discuss three recommendations that you might suggest.Diet Analvsis: Fat (g) Food Calories Reduced Fat Blueberry muffin410 10 490 Large Latte 2 slices Bologna, all beef 174 138 2
Answer #1

1. According to Fred's height, weight, and age he should consume of 2200calories.

2. The calories consumed by him is 4015 calories.

3. Percentage of calories from fat:

Total calories = 4015

Fat consumed = 227.3

calories from fat = 227.3/ 4015 = 0.0566

Percentage of calories from fat = 0.0566 X 100

Percentage of calories from fat = 5.6%

4. Recommended maximum % of calories from fat

Recommended calories = 2200

Recommended Maximum calories from fat = 770

Calories from fat = 770/2200 = 0.35

Percentage of calories from fat = 0.35X 100

Percentage of calories from fat = 35%

5. Serving of each food group from fat

Fruits = 2 cups which include raw, frozen fruit, dried fruit, and fruit juice.

Vegetables = 21/2 cup which includes raw, canned, or cooked vegetables, green leafy vegetables.

Grains = 6 ounces which include cereals, bread, cooked rice, pasta.

Protein = 51/2 ounces consists of poultry products, meat, egg, peanut butter, nuts, and seeds.

Diary = 3 cups which includes milk, yogurt, cheese, and soy.

6.a. Vegetables are inadequate in his diet.

b. Concepts to encourage a healthy eating habit

  • Balanced calories
  • Include more vegetables and fruits.
  • cut of sugar and sodium

c. Messages to Fred

  • Make small changes in your daily diet.
  • Be active yourself.
  • Build positive habits.
  • Make a goal
  • Reward yourself for every positive change.